# Any Image Restoration with Efficient Automatic Degradation Adaptation

### v1

@@ 4 Experiments and Comparisons
We conduct experiments adhering to the protocols of prior general image restoration works [57, 93] under two settings: (i) All-in-One and (ii) Single-task. In the All-in-One setting, a unified model is trained to handle multiple degradation types, considering three and five distinct degradations. In the Single-task setting, separate models are trained for each specific restoration task.
[…]
Datasets. For both All-in-One and single-task settings, we follow the protocols of existing work [33, 57] and include the following datasets: For image denoising in single task setting, we combine the BSD400 [1] and WED [51] datasets, adding Gaussian noise at levels $\sigma\in[15,25,50]$ to create noisy images. Testing is conducted on the BSD68 [52] and Urban100 [26] datasets. For single-task deraining, we use Rain100L [81]. The single task dehazing task utilizes the SOTS [36] dataset. For deblurring and low-light enhancement, we employ the GoPro [55] and the LOL-v1 [78] dataset, respectively. To develop a unified model for all tasks, we merge these datasets in either a three or five degradation setting, training for 130 epochs and directly evaluating across different tasks. For single tasks, our method is trained for 130 epochs on the respective training set.
@@
AnyIR (ours) 6M 29.84 .977 36.91 .977 31.15 .882 26.86 .822 23.50 .845 29.65 .901
@@
AnyIR (ours) 6M 34.27 .937 31.63 .896 28.39 .810
{tblr}
colspec = @lX[c]X[c]X[c]X[c]X[c]@, colsep=0.1pt, rows=rowsep=0.75pt, stretch = 0, & Input AirNet [33] PromptIR [57] Ours GT
[…]
{tblr}
[…]
{tblr}
[…]
Figure 4: Visual comparison of AnyIR with state-of-the-art methods considering three degradations. Zoom in for better view.
@@
Three Degradations. We compare our All-in-One restorer AnyIR with specialized All-in-One restoration methods, including BRDNet [70], LPNet [21], FDGAN [16], DL [20], MPRNet [91], AirNet [33], and PromptIR [57], trained simultaneously on three degradations, i.e., dehazing, deraining, and denoising. Our proposed AnyIR model consistently outperforms other methods as the leading All-in-One image restorer, demonstrating an average improvement of 0.45 dB across all benchmarks. It also stands out as the most efficient architecture, as demonstrated in Tab. 1. Notably, our method achieves state-of-the-art performance on the SOTS and Rain100L benchmarks, surpassing the previous best, PromptIR [57], by 0.38 dB and 1.83 dB, respectively. In terms of denoising, AnyIR achieves the second-highest PSNR in all $\sigma$ settings. Furthermore, it accomplishes this with approximately 80% fewer parameters and FLOPs
@@
Five Degradations. Building on recent studies [33, 93], we extend the three degradation settings to include deblurring and low-light image enhancement, thereby validating our method’s effectiveness in a more comprehensive All-in-One setting. As shown in Tab. 2, our AnyIR model excels by effectively enhancing the extracted features with degradation-specific knowledge. AnyIR outperforms AirNet [33] and IDR [93] by an average of $4.16$ dB and $1.22$ dB, respectively, across all five benchmarks, while using $33\%$ and $60\%$ fewer parameters. Furthermore, we compare our AnyIR model to general image restorers trained under the same All-in-One settings. Notably, our method outperforms Restormer [89] and NAFNet [6] on the LOLv1 dataset by 3.09 dB and 3.01 dB, respectively. Additionally, our model is four times smaller than Restormer and three times smaller than NAFNet.
@@
Single-Degradation. To assess the efficacy of our proposed framework, we present results in Tab. 3 wherein individual instances of our method are trained using the single degradation protocol for dehazing, deraining, and denoising. For dehazing, AnyIR archives the second-best PSNR and the best SSIM performance. For draining, our AnyIR outperforms the previous state-of-the-art method PromotIR [57] by a large margin, i.e., 0.75 dB and 0.003 in terms of the PSNR and the SSIM. For denoising, AnyIR also ranks as the runner-up among all comparison methods, underscoring its effectiveness in reducing noise while preserving image details. In addition, by comparing the single-degradation and multi-degradation performances, we can observe that all other models suffer from significant performance drops, while our model achieves more robust overall performance. This can be attributed to our unified cross-degradation embedding. Finally, our model exhibits a significantly smaller size compared to other approaches, emphasizing the efficiency and practical feasibility of the proposed AnyIR.
Visual results. To complement the quantitative results, we visualize the outcomes of our method in Fig. 4. The visualizations demonstrate the efficacy of AnyIR across different image restoration tasks, including dehazing, deraining, and denoising. In the dehazing task, both AirNet [33] and PromptIR [57] exhibit limitations in fully eliminating haziness, leading to noticeable color reconstruction discrepancies. A typical example can be found in the first-row image where the restoration fails for the Dark Gray building. In contrast, our AnyIR method effectively enhances visibility and ensures precise color reconstruction.
In challenging rainy scenes, previous methods continue to exhibit remnants of rain streaks. Please, zoom in for more details. In contrast, our approach excels at eliminating these artifacts and recovering underlying details, showing our superiority under adverse weather conditions.
Additionally, AnyIR produces clear and sharp denoised outputs, preserving sharp edges and fine textures. This results in cleaner and more visually pleasing images.
Note that the current SOTA PromptIR is approximately 6$\times$ larger than ours. Despite this, our efficient network consistently produces visually superior results, demonstrating its effectiveness. For more detailed examples, please refer to the supplementary material.
@@
AnyIR(ours) 32.51 3556M 6.29M 19G
Tab. 4 compares memory usage, FLOPs, and model parameters, showcasing our framework’s superiority over state-of-the-art all-in-one restorers. By employing a hybrid block design with our proposed degradation adaptation module, AnyIR achieves a PSNR that is 0.45 dB higher than PromptIR [57], while using 63.8% less memory and 82.3% fewer parameters. Most strikingly, AnyIR requires only 19G FLOPs, making it 85.6% more efficient in terms of computational complexity. Our efficiency can be attributed to our design of processing feature modeling in the sub-latent space with our proposed global-local intertwining. Hence, we achieve remarkable efficiency while achieving very competitive performance. We believe that our substantial reductions in memory, parameters, and FLOPs highlight AnyIR’s superior efficiency and practicality for image processing tasks, which have the potential to serve as a fresh baseline for future research.
@@
Figure 5: The structure of different DAB variants.
5 Ablation Studies and Discussions
Ablation Studies. We conduct detailed studies on the proposed components within our AnyIR framework. All experiments are conducted in the All-in-One setting with three degradations. We compare our plain feature modeling block DAB against other variants. As detailed in Tab. 5, we assess the effectiveness of our key architectural contributions by removing or replacing our designed module with the counterparts. The detailed architecture overview of these considered counterparts can be found in Fig. 5.
@@ 5 Ablation Studies and Discussions
Figure 6: The training samples statistics.
Affect of data distribution.
Table 6: The average results for denoising and overall (i.e., dehazing, draining, and denoising) under the 3-degradation IR setting.
Method Denoising Overall
PromptIR [57] 31.11 0.873 32.06 0.913
AnyIR (Original,ours) 31.09 0.873 32.51 0.916
AnyIR (New, ours) 31.24 0.877 32.58 0.919
As illustrated in Fig. 6, the original training samples exhibit significant variation across different IR tasks. Based on this observation, we propose a new set of training samples that over-represent the denoising samples and reduce the dehazing samples by half. This adjustment aims to achieve a more balanced data distribution. The results, presented in Tab. 6, demonstrate that training AnyIR with this revised set of samples, even when using only 10% of the total training samples, yields significant performance improvements.
Discussion. The All-in-One IR is a relatively new research direction compared to degradation-specific ones. Pioneering works aggregated data from various degradation fields to form new training samples, then tested on each degradation dataset. While this approach marked progress, we observe a significant drawback: it overlooks the data distribution across degradation domains. The current training data is unbalanced, with a disproportionate amount of data coming from the dehazing domain, which contributes to 67% of the total samples. In contrast, other tasks such as denoising represent only 12%. This imbalance causes the learning model to prioritize one specific degradation over others, which is supported by our experimental results. As shown in Tab 1, our model performs slightly worse on denoising, while achieving state-of-the-art results on dehazing and draining. In addition, the results in Tab. 6 also indicate that a slightly balanced training set has the potential to improve the All-in-One IR significantly. For future research, we recommend considering a more balanced data proportion to improve overall performance.
@@ 6 Conclusion
We introduce AnyIR, an efficient single-image restoration model capable of handling any type of degradation. Our model leverages the intrinsic characteristics of each degradation as gated guidance, directing the overall feature modeling process. This approach enables the joint embedding of all degradation types while preserving the unique attributes of each. AnyIR sets a new standard in degradation awareness, achieving unparalleled efficiency and fidelity in all-in-one image restoration. We hope that our model will serve as a fresh baseline for future research in this field. In this supplementary material, we first present additional experimental results including the full inference time comparison and additional visual comparison results. Next, we discuss the limitations of our work and potential directions for future research. Finally, we explore the possible broader impacts of our study.
Appendix A Additional Results
Full Inference Time Comparison. Based on the full inference time comparison in Tab. 7, our proposed AnyIR model demonstrates a significant reduction in processing time across all tasks when compared to the PromptIR model. Specifically, for denoising tasks with varying noise levels ($\sigma=15$, $\sigma=25$, and $\sigma=50$), AnyIR consistently achieves a faster inference time of 30 seconds, compared to 41-42 seconds with PromptIR. Similarly, in the deraining task, AnyIR reduces the inference time from 60 seconds to 43 seconds. The most notable improvement is observed in the dehazing task, where AnyIR cuts the inference time from 416 seconds to 274 seconds. These results indicate that our AnyIR model not only maintains high performance but also offers substantial efficiency improvements, making it more suitable for real-time applications on resource-constrained devices
Additional visual results. The visual comparison results under the 3-degradation IR setting shown in Fig. 7 also show that The proposed AnyIR can effectively restore the clean image from its degraded counterparts.
Figure 7: Additional visual comparison under the 3-degradation setting.
Appendix B Limitations and Future Work
Though the proposed AnyIR achieves new state-of-the-art performance under the current all-in-one IR setting, we have observed that the training dataset used for this task is severely unbalanced. This imbalance may lead to unbalanced IR performance. To address this, we plan to work on creating a more balanced training dataset, which will help make the proposed model even more efficient and effective.
Appendix C Broader Impact
The development of our unified image restoration model holds significant potential for a wide range of applications, extending its influence beyond the immediate technical advancements. By dramatically reducing model complexity and computational requirements, our approach can make high-quality image restoration more accessible on mobile devices and other resource-constrained platforms. This accessibility can empower users in various fields such as telemedicine, remote sensing, and digital archiving, where the ability to restore degraded images efficiently and effectively is crucial. Furthermore, by minimizing the computational footprint, our model contributes to reducing the environmental impact associated with large-scale data processing and storage, aligning with sustainable computing practices. The public availability of our code will also encourage further research and development, fostering innovation and collaboration within the scientific community. By setting new standards for efficiency and effectiveness in image restoration, our work not only advances the state-of-the-art but also broadens the scope of practical applications and promotes more sustainable technological growth.
Table 7: Full inference time comparison under the 3-degradation all-in-one IR setting. The lower the better.
Task PromptIR [57] AnyIR (Ours)
Denoising ($\sigma=15$) 41s 30s
Denoising ($\sigma=25$) 41s 30s
Denoising ($\sigma=50$) 42s 30s
Deraining 60s 43s
Drhazing 416s 274s

### v2

@@ 4 Experiments and Comparisons
We conduct experiments adhering to the protocols of prior general image restoration works [53, 88] under two settings: (i) All-in-One and (ii) Single-task. In the All-in-One setting, a unified model is trained to handle multiple degradation types, considering three and five distinct degradations. In the Single-task setting, separate models are trained for each specific restoration task. The implementation details and the dataset introduction are provided in our Supp. Mat.
4.1 State of the Art Comparison
@@
PromptIR [53] 36M 26.54 .949 36.37 .970 31.47 .886 28.71 .881 22.68 .832 29.15 .904
RAM (ours) 6M 29.84 .977 36.91 .977 31.15 .882 26.86 .822 23.50 .845 29.65 .901
Methods with the assistance of natural language prompts and large vision model
InstructIR [11] 16M 36.84 .973 27.10 .956 31.40 .887 29.40 .886 23.00 .836 29.55 .907
Perceive-IR [93] 42M 28.19 .964 37.25 .977 31.44 .887 29.46 .886 22.88 .833 29.84 .909
@@
RAM (ours) 6M 34.27 .937 31.63 .896 28.39 .810
@@
Three Degradations. We evaluate our All-in-One restorer, RAM , against other specialized methods, including BRDNet [66], LPNet [23], FDGAN [17], DL [22], MPRNet [85], AirNet [34], and PromptIR [53], all trained on three degradations: dehazing, deraining, and denoising. RAM consistently leads as the top-performing All-in-One restorer, with an average gain of 0.45 dB across benchmarks and the most efficient architecture, as shown in Tab. 1. It achieves state-of-the-art results on SOTS and Rain100L, outperforming PromptIR by 0.38 dB and 1.83 dB, respectively. In denoising, RAM ranks second across all $\sigma$ levels while maintaining 80% fewer parameters and FLOPs.
@@
Five Degradations. Extending the three degradation tasks to include deblurring and low-light enhancement [34, 88], we validate our method’s comprehensive performance in an All-in-One setting. As shown in Tab. 2, RAM effectively leverages degradation-specific features, surpassing AirNet [34] and IDR [88] by an average of 4.16 dB and 1.22 dB, respectively, with 33% and 60% fewer parameters. Against general image restorers, RAM outperforms Restormer [86] and NAFNet [5] on the LOLv1 dataset by 3.09 dB and 3.01 dB, while being four times smaller than Restormer and three times smaller than NAFNet.
@@
Single-Degradation. We present results in Tab. 3 with RAM trained on individual degradation tasks: dehazing, deraining, and denoising. For dehazing, RAM achieves the second-best PSNR and the highest SSIM. In deraining, RAM surpasses the previous state-of-the-art, PromptIR [53], by 0.75 dB in PSNR and 0.003 in SSIM. For denoising, RAM ranks second overall, effectively reducing noise while preserving details. Notably, while other models show significant performance drops in multi-degradation settings, RAM maintains robust performance, attributed to its unified cross-degradation embedding. Furthermore, RAM is substantially smaller than competing models, highlighting its efficiency and practical feasibility.
Visual results.
Figure 4: Visual comparison of RAM with state-of-the-art methods considering three degradations. Zoom in for a better view.
To complement the quantitative results, we visualize the outcomes of our method in Fig. 4. The visualizations demonstrate the efficacy of RAM across dehazing and draining, and we marked out the detailed region via the red boxes. The denoising visual results are provided in our Supp. Mat. In the dehazing task, both AirNet [34] and PromptIR [53] exhibit limitations in fully eliminating haziness, leading to noticeable color reconstruction discrepancies. In contrast, our RAM method effectively enhances visibility and ensures precise color reconstruction. In challenging rainy scenes, previous methods continue to exhibit remnants of rain streaks. Please, zoom in for more details. In contrast, our approach excels at eliminating these artifacts and recovering underlying details, showing our superiority under adverse weather conditions. Note that the current SOTA PromptIR is approximately 6$\times$ larger than ours. Despite this, our method consistently produces visually superior results, demonstrating its effectiveness. For more detailed examples, please refer to the Supp. Mat.
@@
RAM (ours) 32.51 3556M 6.29M 19G
Tab. 4 compares memory usage, FLOPs, and model parameters, highlighting our framework’s efficiency over state-of-the-art all-in-one restorers. With a hybrid block design and our proposed degradation adaptation module, RAM achieves a 0.45 dB PSNR improvement over PromptIR [53] while using 63.8% less memory, 82.3% fewer parameters, and only 19G FLOPs, making it 85.6% more computationally efficient. This efficiency is due to our global-local intertwining in the sub-latent space, enabling both high performance and practical applicability. Our significant reductions in memory, parameters, and FLOPs position RAM as a strong, efficient baseline for future research.
@@
Figure 5: Structure of different DAB variants.
4.2 Ablation Studies
Impact of different components. We conduct detailed studies on the proposed components within our RAM framework. All experiments are conducted in the All-in-One setting with three degradations. We compare our plain feature modeling block DAB against other variants. As detailed in Tab. 5, we assess the effectiveness of our key architectural contributions by removing or replacing our designed module with the counterparts. The detailed architecture overview of these considered counterparts can be found in Fig. 5.
@@ 4.2 Ablation Studies
Table 6: Results of different ($\alpha$, $\beta$, $\gamma$) for three-degradation.
($\alpha$, $\beta$, $\gamma$) PSNR (dB, $\uparrow$) SSIM ($\downarrow$)
(0, $\nicefrac{{hidden}}{{2}}$, $\nicefrac{{hidden}}{{2}}$) 32.14 .910
($\nicefrac{{hidden}}{{2}}$, 0, $\nicefrac{{hidden}}{{2}}$) 32.21 .912
($\nicefrac{{hidden}}{{2}}$, $\nicefrac{{hidden}}{{2}}$, 0) 31.37 .907
($\nicefrac{{hidden}}{{4}}$, $\nicefrac{{hidden}}{{4}}$, $\nicefrac{{hidden}}{{2}}$) 32.51 .916
Impact of different $\alpha$, $\beta$, and $\gamma$ values in Gated DA. Tab.6 shows the different impact of different value of ($\alpha$, $\beta$, $\gamma$) in the proposed gatedDA. We noticed that when setting one of these parameters to 0, the performance is decreased. When we set ($\alpha$, $\beta$, $\gamma$) to ($\nicefrac{{hidden}}{{4}}$, $\nicefrac{{hidden}}{{4}}$, $\nicefrac{{hidden}}{{2}}$), the average performance is increased. This means each part of the proposed gatedDA is necessary. The visual features shown in Fig. 6 indicate that $\alpha$, $\beta$, and $\gamma$ consistently focus on various aspects of the degraded regions, each specializing in different levels or types of degradation.
Figure 6: Visual feature maps of $\alpha$, $\beta$, and $\gamma$ within GatedDA.
Figure 7: Visulization of the channel activation distribution.
5 Discussion
What does the proposed Skip-Split bring? Adjacent channels often contain redundant information due to spatial correlations in the data. Directly splitting the channels into two contiguous halves can lead to uneven feature distribution, with one half potentially capturing redundant features while the other lacks important information. By using a simple skip-split method that interleaves channels between the two processing paths, we ensure a more balanced and diverse set of features in each path, enhancing the effectiveness of both the self-attention and convolutional components. These phenomena is also validated in Fig. 7.
Why does the proposed GatedDA work? As shown in Fig. 8, the error map visualization reveals that degradation in an input image is often not evenly distributed, instead presenting as both global patterns and localized clusters. This suggests that effective degradation modeling must capture both widespread and specific distortions. The proposed GatedDA module addresses this need by targeting its feature responses specifically on degraded regions, achieving a close alignment with the error map. This focus underscores GatedDA’s capacity to selectively enhance degradation features, and when combined with global attention, it provides a comprehensive representation of the image’s degradation structure. This synergy between GatedDA and global attention improves overall restoration quality by allowing nuanced handling of various degradation types. Importantly, our results demonstrate that GatedDA maintains this adaptability across diverse degradation conditions, emphasizing its robustness and versatility in restoration tasks.
Why is RAM efficient? RAM reduces computational costs by splitting input channels: one half is processed by self-attention, the other by a Gated CNN block. This division halves the self-attention’s complexity from $O(B\cdot\text{head}\cdot(H\cdot W)^{2})$, while the Gated CNN processes the remaining channels with a lower complexity of $O(B\cdot C\cdot H\cdot W)$, especially beneficial for high-resolution inputs where $(H\cdot W)^{2}$ dominates. Additionally, RAM employs dimensionality reduction and parameter-efficient designs, collectively reducing GFLOPs and model parameters. This balance of global context modeling and efficient local feature extraction enables RAM to minimize computational costs without sacrificing performance.
Figure 8: Visulization of the error map and the output of GatedDA.
Is scaling down the new advantage? While recent methods in image restoration often scale up model size and complexity, our approach takes a different path: scaling down. Instead of relying on large architectures, we embrace a simple but non-trivial design, using targeted components like the GatedDA module to effectively capture degradation without excessive parameters. This simplicity is not a compromise; it’s an asset—yielding high-quality restoration with minimal computational demand, faster training, and greater adaptability. The All-in-One IR framework, though new compared to degradation-specific methods, faces a key limitation: an imbalance in data distribution, with certain degradations (e.g., dehazing) dominating. Our experiments suggest that a more balanced dataset could significantly improve performance across degradation types, offering a constructive direction for future work. Additional analyses, experiments, and visual results are available in our Supp. Mat.
@@ 6 Conclusion
We introduce RAM , an efficient single-image restoration model capable of handling any type of degradation. Our model leverages the intrinsic characteristics of each degradation as gated guidance, directing the overall feature modeling process. This approach enables the joint embedding of all degradation types while preserving each of their unique attributes. RAM sets a new standard in degradation awareness, achieving unparalleled efficiency and fidelity in all-in-one image restoration. We hope that our model will serve as a fresh baseline for future research in this field.